OpenStudy (vera_ewing):

TRUE or FALSE? When Point E (-2, -5) is rotated 270° counterclockwise about the origin, it becomes (-5, 2).

OpenStudy (vera_ewing):

@wio

OpenStudy (vera_ewing):

@amistre64 please help me!

OpenStudy (amistre64):

any way to check it that you know of?

OpenStudy (vera_ewing):

well i guess i could do the actual problem? lol

OpenStudy (amistre64):

lol, that would be one method yes. i know of several more

OpenStudy (vera_ewing):

oh really? like what?

OpenStudy (amistre64):

well, since 270 = -90, then the dotproduct of the 2 vectors would equal 0, and have the same length

OpenStudy (amistre64):

another method is a rotation matrix -2cos(270) +5sin(270) = 5 -2sin(270) -5cos(270) = -2

OpenStudy (vera_ewing):

ok so its false?

OpenStudy (amistre64):

did you work your method? or is that just a guess?

OpenStudy (vera_ewing):

well i worked the method with the rotation matrix

OpenStudy (amistre64):

-2,-5 -5, 2 ------ 10-10 = 0 so they are 90 from each other.  im getting false as well this way